Circular No. 8653

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 A1 (POJMANSKI) Grzegorz Pojmanski, Warsaw University Astronomical Observatory,

reports his discovery of a diffuse comet with a coma diameter of 1′ (FWHM) on 3-min CCD frames (pixel size 14′′.8) taken on 2006 Jan. 1 and 4 with a 180-mm-focal-length f/2.8 telephoto lens (+ V filter) in the course of the All Sky Automated Survey at Las Campanas; Pojmanski subsequently identified ASAS images of the comet taken on 2005 Dec. 29. S. Roland (Montevideo, Uruguay) writes that 20-second CCD exposures taken with a 0.46-m prime-focus telescope confirm the cometary nature of the new object, with a coma diameter of ∼ 1′. More than seven hours after Pojmanski reported his find to the Central Bureau, K. Cernis reported (apparently independently) that he has found an object on ultraviolet SWAN/SOHO website images taken on three nights (back to 2005 Dec. 25) showing a motion consistent with that of C/2006 A1.

2005/6 UT  α_{2000}  δ_{2000}  Mag. Observer
Dec. 25  22^{h}09.^{m}0  −71^{o}07′  SWAN
29  21 52.6  −69 03  _{”}
29.03145 21 50 29^{s}  −69 40.6  13.2 ASAS
Jan. 1  21 42.4  −68 23  SWAN
1.04388 21 40 28  −68 36.5  12.5 ASAS
4.03225 21 31 33  −67 31.2  12.0  _{”}
4.03734 21 31 33  −67 31.1  12.0  _{”}
4.99875 21 28 51.64 −67 09 39′′.9 14.6 Roland
5.00339 21 28 50.91 −67 09 33.7 14.6  _{”}
5.01031 21 28 49.73 −67 09 23.9 14.7  _{”}
5.01681 21 28 48.72 −67 09 15.0 14.6  _{”}
5.02197 21 28 47.83 −67 09 07.9 14.8  _{”}
5.03106 21 28 46.33 −67 08 55.5 14.5  _{”}
5.03985 21 28 44.88 −67 08 43.5 14.7  _{”}
5.04632 21 28 43.71 −67 08 35.7 14.2  _{”}
5.05587 21 28 42.25 −67 08 21.4 14.7  _{”}
5.07266 21 28 39.51 −67 07 58.5 14.5  _{”}

Preliminary parabolic orbital elements from the ASAS and Roland data:

T = 2006 Feb. 22.671 TT  ω = 350.^{o}499
Ω = 212.275 }2000.0
q = 0.56748 AU  i = 93.230
2006 January 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green